OpenStudy (precal):

I need to find the tangent line of f(x)=e^(3x) any help is greatly appreciated

OpenStudy (precal):

I know I need to take the derivative

OpenStudy (anonymous):

so the derivative of e^(ax) is ae^(ax)

OpenStudy (precal):

\[f '=3e^{3x}\]

OpenStudy (anonymous):

correct

OpenStudy (precal):

yes and then I sub it into y-y1=m(x-x1) sorry I should use subscripts

OpenStudy (anonymous):

yes you evaluate the derivative at the requested point and it gives you the slope m

OpenStudy (anonymous):

then you use the coordinates of the point as x1 and y1

OpenStudy (precal):

that is where I am messing up

OpenStudy (precal):

\[y-y_{1}=m(x _{1}-x)\]

OpenStudy (anonymous):

no, y-y1=m(x-x1)

OpenStudy (precal):

\[y-1=3e^{3x}(x-0)\]

OpenStudy (anonymous):

"you evaluate the derivative at the requested point and it gives you the slope m" you need to plug into the derivative in your case x=0 and then plug it as m.

OpenStudy (anonymous):

m = f'(x1)

OpenStudy (precal):

so m=3

OpenStudy (precal):

now I use the slope intercept correct? it has been a long time since I have seen this.....

OpenStudy (precal):

ok got it y=3x+1 you are a lifesaver.....

OpenStudy (anonymous):

good !

OpenStudy (precal):

Thank you